# Category-level Meta-learned NeRF Priors for Efficient Object Mapping

### v2

@@ Abstract
In 3D object mapping, category-level priors enable efficient object reconstruction and canonical pose estimation, requiring only a single prior per semantic category (e.g., chair, book, laptop). Recently, DeepSDF has predominantly been used as a category-level shape prior, but it struggles to reconstruct sharp geometry and is computationally expensive. In contrast, NeRFs capture fine details but have yet to be effectively integrated with category-level priors in a real-time multi-object mapping framework. To bridge this gap, we introduce PRENOM, a Prior-based Efficient Neural Object Mapper that integrates category-level priors with object-level NeRFs to enhance reconstruction efficiency while enabling canonical object pose estimation. PRENOM gets to know objects on a first-name basis by meta-learning on synthetic reconstruction tasks generated from open-source shape datasets. To account for object category variations, it employs a multi-objective genetic algorithm to optimize the NeRF architecture for each category, balancing reconstruction quality and training time. Additionally, prior-based probabilistic ray sampling directs sampling toward expected object regions, accelerating convergence and improving reconstruction quality under constrained resources. Experimental results on a low-end GPU highlight the ability of PRENOM to achieve high-quality reconstructions while maintaining computational feasibility. Specifically, comparisons with prior-free NeRF-based approaches on a synthetic dataset show a 21% lower Chamfer distance, demonstrating better reconstruction quality. Furthermore, evaluations against other approaches using shape priors on a noisy real-world dataset indicate a 13% improvement averaged across all reconstruction metrics, and comparable pose and size estimation accuracy, while being trained for 5× less time.
@@ I Introduction
These priors can be broadly classified into three categories. Instance-level priors [8, 9] offer high accuracy but require knowledge of the environment, through precise CAD models of each object instance, limiting their applicability to specialized use cases. On the other extreme, generalized priors [10] use a single prior for all objects but rely on bulky diffusion models, making them computationally intensive even for a single object. Category-level priors are a suitable trade-off as they rely solely on semantic labels, and thus can work with unseen objects within predefined categories of interest. Recent works represent such priors using variational autoencoders [11], shape embeddings [12], and DeepSDFs [13, 14, 15]. While these methods offer reasonable global shape and pose estimation, they are computationally demanding, requiring high-end GPUs; use identical prior designs for all categories, ignoring differences in shape complexity; and in most scenarios struggle with reconstructing sharp geometry.
To address these limitations, we present PRENOM, a Prior-based Efficient Neural Object Mapper that represents objects by lightweight NeRF models, with meta-learned [16] parameter initializations as category-level priors, within a real-time RGB-D multi-object mapping framework. Having been trained on large shape datasets, these priors can localize objects within a canonical frame while also benefiting from the high-fidelity reconstructions typically associated with NeRFs. Further integrating them with neural architectures optimized for each category and prior-based probabilistic ray sampling, we achieve faster training convergence (see progress comparison in Figure 1), enabling accurate reconstructions, even with limited GPU resources.
@@ II-B Category-level Priors for Object Mapping
NodeSLAM [11] employs variational autoencoders as category-level shape priors to predict occupancy grids, jointly optimizing the shape and pose of objects. However, it struggles with reconstructing complex shapes. Frodo [12] represents shape priors using encoder-decoder networks to model objects as point clouds and SDFs. Unlike our sequential method, it operates in batch mode.
Several works [13, 14, 15] adopt the DeepSDF architecture [19] as a shape prior, learning latent embeddings as shape priors, and training decoder networks to predict object SDFs. These approaches require long training times of several seconds per object on a low-end GPU while having limited capability to reconstruct fine geometry. Moreover, these methods also jointly optimize the object pose with shape reconstruction. While this can refine object localization, it increases training time and makes reconstruction accuracy dependent on state estimation quality. In contrast, our approach follows [11] by estimating the object pose, before initializing reconstruction. However, unlike the CNNs used by [11], we rely solely on our priors to estimate and refine the pose.
@@ III-A Offline Prior Training
where $\eta$ is the task learning rate. The loss function is based on differentiable volume rendering described later in Section III-B. After a set number of inner iterations $q$, the updated parameters $\theta_{\mathbf{i}}^{\,(q)}$ are used to update the meta-parameters via a simple interpolation:
@@
where $\beta$ is the meta-learning step size. By meta-learning on several reconstruction tasks, the prior learns common geometric features within a category, an example of which can be seen in Figure 3.
Architecture Optimization To choose $\hat{\alpha}$, we perform a multi-objective optimization, using a mixed-variable genetic algorithm [25] with the objectives of minimizing the training time and the object reconstruction error (measured using Chamfer distance). At each optimization iteration, meta-learning is used to obtain the optimized parameters $\theta_{\mathbf{k},\mathbf{N}}$, as described in Equations 1, LABEL: and 2. These learned parameters are then evaluated on $\mathcal{T}^{test}_{\mathbf{k}}$ to guide the genetic algorithm to generate an improved architecture for the next iteration. Starting with an initial population of architectures, this iterative process is repeated for $\mathbf{M}$ generations, producing a set of Pareto-optimal solutions. From these solutions, the final architecture $\hat{\alpha}_{\mathbf{k}}$ is chosen as the knee point of the convex decreasing curve that best fits the solution set (see Figure 4). Consequently, $\hat{\theta}_{\mathbf{k}}$ are the corresponding meta-learned parameters for the selected architecture.
2.22.22.22.22.32.32.32.32.42.42.42.42.52.52.52.52.62.62.62.62.72.72.72.75.385.385.385.385.405.405.405.405.425.425.425.425.445.445.445.445.465.465.465.46Chosen architectureChamfer distance (mm)Scaled training time per iteration (s)Possible architecture for $\mathbf{BOOK}$
@@ III-B Online Object Mapping
Coarse Object State Estimation The 9-DoF state comprises a 6-DoF pose and a 3-DoF size. Consistent with related works, we assume that objects are parallel to the ground, reducing the pose estimation problem to position and yaw estimation. Following [6, 1], the optimal yaw is determined by aligning line features extracted from images with the edges of the bounding box. Once yaw is estimated, the object size is computed from the lengths of its spatial extents, while its position is defined as the center of these extents.
@@
An object is considered ready for reconstruction after it has been observed from a sufficiently wide range of viewpoints. If the object has a prior (semantic class $\mathbf{k}\in\mathcal{I}$), its pose is transformed into the canonical frame relative to which the prior was learned. For this, we sample yaw angles between 00 and $2\pi$, and for each sample, apply the corresponding yaw rotation to the object point cloud $P$ to obtain $P_{\gamma}$. The object’s size $\mathcal{S}_{\gamma}$ is estimated from the spatial extents of $P_{\gamma}$ and is used to normalize the point cloud, producing $\overline{P}_{\gamma}=P_{\gamma}/\mathcal{S}_{\gamma}-1/2$. The optimal yaw angle $\hat{\gamma}$ is then selected by maximizing the accumulated density of the normalized points in the prior density grid:
@@
Object-level NeRFs Once the object state is estimated, it is implicitly reconstructed using a lightweight NeRF model, which incorporates a multi-resolution hash encoding [28] alongside a shallow MLP. Supervision is achieved through differentiable volume rendering, which samples and accumulates color and opacity along camera rays to reconstruct the object. For a known object $\mathbf{k}\in\mathcal{I}$, the optimized initial parameters $\hat{\theta}_{\mathbf{k}}$ are loaded into a NeRF model with architecture $\hat{\alpha}_{\mathbf{k}}$. Otherwise, a default architecture (adapted from [6]) and random parameter initialization are used (refer to Figure 2 for color-coded flow). All object-level NeRFs are trained in parallel using multi-threading.
Training Data RGB-D images, instance masks, and camera poses at every keyframe are sent to the GPU as training data.
@@
We implement this sampling using CUDA kernels, ensuring minimal computational overhead.
@@ IV Evaluation
IV-A Experimental Setup
To train piors, we run $\mathbf{M}=100$ generations of the genetic algorithm to optimize the architectures $\{\alpha_{\mathbf{k}}\,|\,\mathbf{k}\in\mathcal{I}\}$. Each $\alpha$ consists of all tunable hyperparameters of the NeRF representation (refer to [28] for a full list), the meta-learning ($\mathbf{N}$, $q$, $\eta$, and $\beta$), and the volume rendering loss ($\lambda_{\mathbf{d}}$ and $\lambda_{\sigma}$). For object reconstruction during online mapping, 4096 rays are generated at each iteration with $N_{c}=N_{r}=32$ samples per ray. Moreover, we set $R=64$, $\epsilon=10^{-4}$ and $m=50$.
We run the experiments on a system equipped with an Intel Core i9-12900H and a low-power, low-memory (4GB) NVIDIA T600 Laptop GPU. This setup is ideal for evaluating the efficiency of our approach, as only a limited number of training iterations per object is possible to maintain sequential real-time processing of multi-object scenarios.
@@ IV-B Comparison with prior-free methods
We assess the effectiveness of our priors by comparing against prior-free methods that also use NeRFs as the object representation.
@@ IV-C Comparison with category-level shape prior approaches
For object state estimation, we compute translation, rotation, and size errors. Morover, similar to PRENOM, the baseline methods also assume that chairs lie parallel to the ground to compute the initial pose. However, since these methods update the pose during reconstruction, we align their final poses’ vertical axes with the ground truth before evaluation to ensure fairness.
Results and discussion PRENOM achieves an improvement in reconstruction quality of approximately 13% averaged across all metrics compared to the second-best approach, as shown in Table II. Notably, it outperforms other methods in reconstruction quality while only being trained for 5$\times$ less time and with a 30% smaller prior size, underscoring its efficiency. Additionally, comparisons with our prior-free version re-emphasize that our prior-based approach not only enables canonical object state estimation but also significantly enhance reconstruction quality.
@@
Furthermore, example qualitative results in Figure 9 demonstrate PRENOM’s ability to reconstruct sharp geometrical features within the observed regions. However, we do not impose geometric constraints with the initial prior, and during training, scene information gradually dominates, leading to noisier reconstructions in partially-observed/unobserved regions and regions with occlusions. In contrast, DeepSDF-based methods struggle to reconstruct complex and thin geometries typical of chairs but can predict the unseen regions in a way that remains structurally similar to a chair, even if it is not consistent with the scene. Future work will focus on incorporating constraints with our prior to mitigate the forgetting issue.
@@
Moreover, despite not jointly optimizing object state with reconstruction, our method achieves comparable object state recovery results as can be seen in Table III. We do not consider the improvement in rotation accuracy to be significant, as PRENOM does not adjust pose during reconstruction and thus benefits from the initial assumption that objects are parallel to the ground. While this initial assumption is applied to all methods, the baseline approaches are more susceptible to noise as they optimize pose during the reconstruction process, which negatively impacts the rotation accuracy.
@@ V Conclusion
We presented PRENOM, an efficient prior-based neural object mapper that leverages category-level priors to improve 3D object reconstruction and enables canonical pose estimation. By integrating meta-learned priors and probabilistic ray sampling, optimized per-category using a multi-objective genetic algorithm, PRENOM enhances reconstruction accuracy and efficiency, making it well suited for multi-object mapping in resource-constrained setups. Experimental results demonstrate its effectiveness, achieving a 21% lower Chamfer distance compared to prior-free NeRF-based representations. Further comparisons with DeepSDF-based approaches show a 13% improvement in reconstruction metrics, while maintaining comparable pose and size estimation capabilities, highlighting the efficiency of our approach, as it was trained for 5× less time and with a 30% smaller prior size.
Despite its advantages, PRENOM is inherently limited by its NeRF-based representation, which struggles with out-of-view feature reconstruction. Although our priors help mitigate these issues, the absence of hard constraints to enforce consistency eventually leads to noisier reconstructions, especially if trained for extended periods. Additionally, object state is not refined during training, resulting in slightly worse state estimation than some existing shape prior methods. Future work will focus in these directions.

### v3

@@ Abstract
In 3D object mapping, category-level priors enable efficient object reconstruction and canonical pose estimation, requiring only a single prior per semantic category (e.g., chair, book, laptop, etc.). DeepSDF has been used predominantly as a category-level shape prior, but it struggles to reconstruct sharp geometry and is computationally expensive. In contrast, NeRFs capture fine details but have yet to be effectively integrated with category-level priors in a real-time multi-object mapping framework. To bridge this gap, we introduce PRENOM, a Prior-based Efficient Neural Object Mapper that integrates category-level priors with object-level NeRFs to enhance reconstruction efficiency and enable canonical object pose estimation. PRENOM gets to know objects on a first-name basis by meta-learning on synthetic reconstruction tasks generated from open-source shape datasets. To account for object category variations, it employs a multi-objective genetic algorithm to optimize the NeRF architecture for each category, balancing reconstruction quality and training time. Additionally, prior-based probabilistic ray sampling directs sampling toward expected object regions, accelerating convergence and improving reconstruction quality under constrained resources. Experimental results highlight the ability of PRENOM to achieve high-quality reconstructions while maintaining computational feasibility. Specifically, comparisons with prior-free NeRF-based approaches on a synthetic dataset show a 21% lower Chamfer distance. Furthermore, evaluations against other approaches using shape priors on a noisy real-world dataset indicate a 13% improvement averaged across all reconstruction metrics, and comparable pose and size estimation accuracy, while being trained for 5$\times$ less time.
Code available at: https://github.com/snt-arg/PRENOM
@@ I Introduction
These priors can be broadly classified into three categories. Instance-level priors [8, 9] offer high accuracy, but require knowledge of the environment, through precise CAD models of each object instance, limiting their applicability to specialized use cases. On the other end, generalized priors [10] use a single prior for all objects but rely on bulky diffusion models, making them computationally intensive even for a single object. Category-level priors are a suitable trade-off as they rely solely on semantic labels, and thus can work with unseen objects within predefined categories of interest. Recent works represent such priors using variational autoencoders [11], shape embeddings [12], and DeepSDFs [13, 14, 15]. While these methods offer reasonable global shape and pose estimation, they are computationally demanding, requiring high-end GPUs; use identical prior designs for all categories, ignoring differences in shape complexity; and in most scenarios struggle with reconstructing sharp geometry.
To address these limitations, we present PRENOM, a Prior-based Efficient Neural Object Mapper that represents objects by lightweight NeRF models, with meta-learned [16] parameter initializations as category-level priors, within a real-time RGB-D multi-object mapping framework. Having been trained on large shape datasets, these priors can localize objects within a canonical frame while also benefiting from the high-fidelity reconstructions typically associated with NeRFs. Further integrating them with neural architectures optimized per-category and prior-based probabilistic ray sampling, we achieve faster training convergence (see Figure 1), enabling accurate reconstructions with limited GPU resources.
@@ II-B Category-level Priors for Object Mapping
NodeSLAM [11] employs variational autoencoders as category-level shape priors to predict occupancy grids, jointly estimate the shape and pose of objects. However, it struggles with reconstructing complex shapes. Frodo [12] represents shape priors using encoder-decoder networks to model objects as point clouds and SDFs. Unlike our sequential method, it operates in batch mode.
Several works [13, 14, 15] adopt the DeepSDF architecture [19] as a shape prior, learning latent embeddings as shape priors, and training decoder networks to predict object SDFs. These approaches require long training times of several seconds per object on a low-end GPU while having limited capability to reconstruct fine geometry. Moreover, these methods also jointly estimating the object pose with shape reconstruction. While this can refine object localization, it increases training time and makes reconstruction accuracy dependent on state estimation quality. In contrast, our approach follows [11] by estimating the object pose, before initializing reconstruction. However, unlike the CNNs used by [11], we rely solely on our priors to estimate and refine the pose.
@@ III-A Offline Prior Training
where $\eta$ is the task learning rate. The loss function $\mathcal{L}_{\mathcal{T}}$ is based on differentiable volume rendering described later in Section III-B. After a set number of inner iterations $q$, the updated parameters $\theta_{\mathbf{i}}^{\,(q)}$ are used to update the meta-parameters via a simple interpolation:
@@
where $\beta$ is the meta-learning step size. By meta-learning on several reconstruction tasks, the prior learns common geometric features within a category, as seen in Figure 3.
Architecture Optimization To choose $\hat{\alpha}$, we perform a multi-objective optimization, using a mixed-variable genetic algorithm [25] with the objectives of minimizing the training time and the object reconstruction error (measured using Chamfer distance). At each optimization iteration, meta-learning is used to obtain the optimized parameters $\theta_{\mathbf{k},\mathbf{N}}$, as described in Equations 1, LABEL: and 2. These learned parameters are then evaluated on $\mathcal{T}^{test}_{\mathbf{k}}$ to guide the genetic algorithm to generate an improved architecture for the next iteration. Starting with an initial population of architectures with size $\mathbf{P}$, this iterative process is repeated for $\mathbf{M}$ generations, producing a set of Pareto-optimal solutions. From these solutions, the final architecture $\hat{\alpha}_{\mathbf{k}}$ is chosen as the knee point of the convex decreasing curve that best fits the solution set (see Figure 4). Consequently, $\hat{\theta}_{\mathbf{k}}$ are the corresponding meta-learned parameters for the selected architecture.
2.22.22.22.32.32.32.42.42.42.52.52.52.62.62.62.72.72.75.385.385.385.405.405.405.425.425.425.445.445.445.465.465.46Chosen architectureChamfer distance (mm)Scaled training time per iteration (s)Possible architecture for $\mathbf{BOOK}$
@@ III-B Online Object Mapping
Coarse Object State Estimation The 9-DoF state comprises a 6-DoF pose and a 3-DoF size. Consistent with related works [6, 1], we assume that objects are parallel to the ground, reducing the pose estimation problem to position and yaw estimation. The optimal yaw is determined by aligning line features extracted from images with the edges of the bounding box. Once yaw is estimated, the object size is computed from the lengths of its spatial extents, while its position is defined as the center of these extents.
@@
An object is considered ready for reconstruction after it has been observed from a predefined sufficiently wide range of viewpoints. If the object has a prior (semantic class $\mathbf{k}\in\mathcal{I}$), its pose is transformed into the canonical frame relative to which the prior was learned. For this, we sample yaw angles between 0 and $2\pi$, and for each sample, apply the corresponding yaw rotation to the object point cloud $P$ to obtain $P_{\gamma}$. The object’s size $\mathcal{S}_{\gamma}$ is estimated as the spatial extents of $P_{\gamma}$ and is used to normalize the point cloud, producing $\overline{P}_{\gamma}=P_{\gamma}/\mathcal{S}_{\gamma}+1/2$ leading to $\overline{P}_{\gamma}\in[0,1]$. The optimal yaw angle $\hat{\gamma}$ is then selected by maximizing the accumulated density of the normalized points in the prior density grid:
@@
Object-level NeRFs Once the object state is estimated, it is implicitly reconstructed using a lightweight NeRF model, which incorporates a multi-resolution hash encoding [28] alongside a shallow MLP. Supervision is achieved through differentiable volume rendering, which samples and accumulates color and opacity along camera rays to reconstruct the object. For a known object $\mathbf{k}\in\mathcal{I}$, the optimized initial parameters $\hat{\theta}_{\mathbf{k}}$ are loaded into a NeRF model with architecture $\hat{\alpha}_{\mathbf{k}}$. Otherwise, a default architecture (adapted from [6]) and random parameter initialization are used. Refer to Figure 2 for color-coded flow.
Training Data RGB-D images, instance masks, and camera poses at every keyframe are used to train the NeRFs.
@@ IV Evaluation
IV-A Experimental Setup and Implementation Details
To train piors, we run the genetic algorithm, to optimize the architectures $\{\alpha_{\mathbf{k}}\,|\,\mathbf{k}\in\mathcal{I}\}$, with initial population size $\mathbf{P}\mathbin{=}25$ and for $\mathbf{M}\mathbin{=}100$ generations. Each $\alpha$ consists of all tunable hyperparameters of the NeRF representation (refer to [28] for a full list), the meta-learning ($\mathbf{N}$, $q$, $\eta$, and $\beta$), and the volume rendering loss ($\lambda_{\mathbf{d}}$ and $\lambda_{\sigma}$). For object reconstruction during online mapping, 4096 rays are generated at each iteration with $N_{c}\mathbin{=}N_{r}\mathbin{=}32$ samples per ray. Moreover, we set $R\mathbin{=}64$, $\epsilon\mathbin{=}10^{-4}$ and $m\mathbin{=}50$. The probabilistic ray sampling is implemented using CUDA kernels ensuring minimal computational overhead. Moreover, a new thread is created for the training of each object-level NeRF.
We run the experiments on a system equipped with an Intel Core i9-12900H and a low-power, low-memory (4GB) NVIDIA T600 Laptop GPU. With this setup, only a limited number of training iterations per object is possible to maintain sequential real-time processing of multi-object scenarios.
@@ IV-B Comparison with prior-free methods
We assess effectiveness by comparing against prior-free methods that also use NeRFs as the object representation.
@@ IV-C Comparison with category-level shape prior approaches
For object state estimation, we compute translation, rotation, and size errors. Morover, similar to PRENOM, the baseline methods also assume that chairs lie parallel to the ground to compute the initial pose. However, since these methods update the pose during reconstruction, we align their final poses’ vertical rotation axes with the ground truth before evaluation to ensure fairness. In contrast, PRENOM only estimates the pose using ¯$\overline{\mathcal{F}}_{\mathbf{CHAIR}}$ and ¯$\overline{\mathcal{M}}_{\mathbf{CHAIR}}$ (Equation 3 for reference) once before reconstructing, so it’s vertical rotation axis is already aligned.
Results and discussion PRENOM achieves an improvement in reconstruction quality of approximately 13% averaged across all metrics compared to the second-best approach, as shown in Table II. Notably, it outperforms other methods in reconstruction quality while only being trained for 5$\times$ less time and with a 30% smaller prior size. Additionally, comparisons with our prior-free version re-emphasize that our prior-based approach not only enables canonical object state estimation but also significantly enhance reconstruction quality.
@@
Furthermore, example qualitative results in Figure 9 demonstrate PRENOM’s ability to reconstruct sharp geometrical features within the observed regions. However, we do not impose geometric constraints with the initial prior, and during training, scene information gradually dominates, leading to noisier reconstructions in partially-observed/unobserved regions and regions with occlusions. In contrast, DeepSDF-based methods struggle to reconstruct complex and thin geometries typical of chairs but can predict the unseen regions in a way that remains structurally similar to a chair, even if it is not consistent with the scene.
@@
Moreover, despite not jointly optimizing object state with reconstruction, our method achieves comparable object state recovery results as can be seen in Table III. We do not consider the improvement in rotation accuracy to be significant, as PRENOM does not adjust pose during reconstruction and thus benefits from the initial assumption that objects are parallel to the ground. While this initial assumption is used by all methods, the baseline approaches are more susceptible to noise as they alter the pose during the optimization process, which can negatively impact the rotation accuracy.
@@ V Conclusion
We presented PRENOM, an efficient prior-based neural object mapper that leverages category-level priors to improve 3D object reconstruction and enables canonical pose estimation. By integrating meta-learned priors and probabilistic ray sampling, optimized per-category using a multi-objective genetic algorithm, PRENOM enhances reconstruction accuracy and efficiency, making it well-suited for multi-object mapping in resource-constrained setups. Experimental results demonstrate its effectiveness, achieving a 21% lower Chamfer distance compared to prior-free obejct-level NeRF-based approaches. Further comparisons with DeepSDF-based approaches show a 13% improvement in reconstruction metrics, with comparable pose and size estimation capabilities, highlighting the efficiency of our approach, as it was trained for 5$\times$ less time and with a 30% smaller prior size. However, despite its advantages, PRENOM is inherently limited by its NeRF-based representation, which struggles with out-of-view feature reconstruction. Although our priors help mitigate these issues, the absence of hard constraints to enforce consistency eventually leads to noisier reconstructions, especially if trained for extended periods. Additionally, object state is not refined during training, resulting in slightly worse state estimation than some existing shape prior methods. Future work will focus in these directions.